# Unification of the Gauge Theories

### v2

@@ Abstract
We treat the Christoffel coefficients as operators and introduce new mappings for quaternionic products to connect with the theory of electrodynamics in general spacetime. By utilizing the directional operator of the covariant derivative, we generalize the quaternionic mechanism to the theory of electrodynamics. We demonstrate that the Einstein equation allows selecting a constant term that aligns with the covariant derivative.
@@ I Introduction
After general relativity successfully explained the universe, many attempts were made to develop a geometric approach to electrodynamics. Unfortunately, most of these efforts were unsuccessful overduin ; ruiz ; wes ; cabral ; goenner ; lobo ; wess ; gross ; ponce ; williams ; wesso . Our current goal is to pursue this objective in a significantly different manner. We aim to incorporate the theory of gravity into the framework of electrodynamics. This integration may enhance our understanding of quantum gravity, a concept that physicists widely seek to realize. To validate the results of this work, we can analyze the components of Riemann’s curvature tensor. The pursuit of quantum gravity as a significant objective has led us to propose a modification in the dimensions of spacetime herbert ; kiefer ; gianluca ; gianlu ; gianluc ; carlo ; ashteka ; ashtekar . In addressing this challenge, we encounter local issues related to quaternionic production, particularly regarding the contraction of indices in tangent and cotangent spaces. However, the fundamental outcome remains consistent, regardless of the chosen space. These new perspectives will be instrumental in advancing our understanding of gravity theory. We examine two structures: the spacetime metric and the quaternionic metric, which involves quaternionic production under a new metric. As will be demonstrated, it is possible to transition from electrodynamics to gravity using quaternions. However, there is a fundamental difference between the two theories. Electrodynamics is a comprehensive theory consisting of two components, described by the strength tensor in four-dimensional space. When extending electrodynamics to a gravity theory using quaternions, we must consider four components in a four-dimensional spacetime. This leads to the idea of component separation. The alteration in the structure of these components results in equations of motion, ultimately culminating in Bianchi’s identity weber ; misner ; wald ; weinberg ; invarno ; malcolm . We will show that space (geometry) has a physical entity. Therefore, equations of motion will have to be determined, which go beyond mathematical relations to the properties of the geometry of space. When space takes on a physical meaning, it must obey the laws of physics, and ”space characters behavior must be affected by the equations of motion”. One of the consequences of this is that Einstein’s equation of gravity is just one of the determining equations of geometry.
@@
Clarification- It is important to clarify definitions and designations: (1) The vector and dual spaces, represented by $\varepsilon$ and $\varepsilon^{\star}$, are ordered and denoted as $\varepsilon^{\star}\otimes\varepsilon$. Other arrangements are not permitted. The notation $\mathsf{u}^{\alpha\beta\cdots}_{\mu\nu\cdots}=u^{\alpha\beta\cdots}_{\mu\nu\cdots}\,(\hat{e}^{\mu}\hat{e}^{\nu}\cdots)\otimes(\hat{e}_{\alpha}\hat{e}_{\beta}\cdots)$ indicates a multicomponent member. The notation is revised as follows: the components $\mathsf{u}^{\alpha\beta}_{\mu\nu}$ will now be expressed as $u^{\alpha\beta}_{\mu\nu}$. The interaction of indices occurs only from a vector space to the dual space and vice versa; in any case, it is from a vector space into the dual space.
@@
The symbol "$\mathbf{g}"$ is the general metric element and comprises $\varepsilon\times\varepsilon$ or $\varepsilon^{\star}\times\varepsilon^{\star}$. Equations 1 and 2 show that the metric (with a repeated index) does not obey the rules of the operator, and it is clear that any other interpretation of equations 1 and 2 would be wrong for the action of the metric. It can be demonstrated that the metric effect does not influence the calculation of the components. Following the definition of "$\mathsf{ast}"$ mapping, the operator $\hat{\Gamma}_{\kappa}$ is represented as follows;
@@
The above statements and results become easier to understand with the new representation because the $\hat{\Gamma}_{\zeta}$ operator yields "$\hat{\Gamma}_{\kappa}=\Sigma_{\alpha,\beta}\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|\equiv\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|"$. The "$\mathsf{ast}"$ mapping becomes more formal when referring to the ”bra- ket” representations. Moreover, the representation corresponding to the ”bra-ket” extends our possibilities for future calculations. The $\alpha$ and $\beta$ are dummy indices in the $\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|$-operator. They are used, among other things, for the gamma operator effect;
@@
In this process, the box is mated with the alpha vector. That is, the box takes the alpha index. Hence,
@@
The relationships between the quaternion generators $\hat{q}_{1}=I$, $\hat{q}_{2}=J$, and $\hat{q}_{3}=K$ are defined within a Euclidean metric and adhere to the following properties: $\hat{q}_{i}\cdot\hat{q}_{j}=\delta_{ij}$ and $\hat{q}_{i}\times\hat{q}_{j}=\Sigma_{k}\epsilon_{ijk}\hat{q}_{k}$. If we include $\hat{q}_{0}=\Im$ (as the identity), we obtain a quaternion represented by the Lorentz quantities (parameters of the Lorentz index). These are coupled and are termed “$q$ vectors.” Lorentz quantities, which carry the Lorentz index, are represented using quaternions. We have two representations of the Lorentz index: the four-vector and the q-vector. We will use the notation $\mathtt{A},\ \mathtt{B},\cdots$ to denote the quaternions. Quaternions are typically constructed from four-vectors, denoted as $A_{\mu}=(A_{0},\mathbf{A})$. They can be expressed as $\mathtt{A}=iA_{0}\Im+A_{k}\hat{q}^{k}\equiv(\imath\mathtt{A}_{0},\mathtt{A})_{q}$, where the $q$-variable signifies the expression $(\imath\mathtt{A}_{0},\mathtt{A}_{k}\hat{q}^{k})$. Bold letters denote the spatial components of the four vectors. The four-vector framework is based on the spacetime metric, while the Euclidean metric pertains to the quaternion representation. The configuration significantly affects quaternionic products. We also analyze the Lorentz index, which connects physical parameters to the quaternion generator. We can generalize the range of coupled quantities from vectors to tensors. For the $q$-vectors $\hat{\Gamma}$ and $\mathtt{C}$, the extended Grassman quaternion multiplication """"$\triangleright$"" ▷ " (with $\ast$-product), is defined as follows andre ; walker ; kansu ; stefano
@@
where "$e"$ is the coupling coefficient constant with a value of one. As mentioned earlier, $\hat{\Gamma}_{\mu}$ contains two coupling indices. Due to the operators $\hat{\Gamma}$’s and the $\ast$-product $\Delta_{\mu}$ is a semi-complete operator. Equation 7 with the $\ast$-product has the expected effects
@@
The covariant derivative is not complete. The incompleteness of the derivative operator is evident from equations I and 5.
$\mu$ and $\zeta$ are free indices in $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$. This means that $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$ ends up being a second rank tensor; In other words, $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}\propto\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}$ actually has the two free indices $\mu$ and $\zeta$ that can be considered as free indices; $\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\equiv\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\ |\mu,\zeta>$. Equation 3, and our calculations in the ”bra-ket” notation show that,
@@
The free indices after transposition can be seen. In this way, $(\Gamma^{\zeta}_{\mu\lambda}\Gamma^{\kappa}_{\nu\zeta}-\Gamma^{\zeta}_{\nu\lambda}\Gamma^{\kappa}_{\mu\zeta})(|\kappa><\lambda|)$ as an operator contains the two free indices $\kappa$ and $\lambda$. It becomes a simplicial $[\hat{\Gamma}_{\mu},\hat{\Gamma}_{\nu}]^{\kappa}_{\ast\ \lambda}$ which has free $\lambda$ and $\kappa$ indices. Next, we follow the computation of $[\Delta_{\alpha},\Delta_{\beta}]_{\ast}$. We compute the case without torsion,
@@ II Electrodynamics
In this section, we present a mechanism whereby the components of gauge fields $\mathbf{A}$ are coupled with the generators of the quaternions ($\mathbf{A}\rightarrow\mathtt{A}$). A connection between the components of the covariant derivative and the quaternion generators is similarly established. The quaternion variables obtained in this way help us to reconstruct the strength tensor of the fields. Due to the semi-completeness of the covariant derivative operator, we will introduce a new quaternion commutator to obtain the equation of motion and conclude the theory.
@@
We show that Maxwell’s equations are accessible with augmented quaternions. The given content confirms the above definitions of the $q$-vectors. We emphasize that the quaternion mapping creates two separate spaces: the tangent and the cotangent $q$-vectors. The measures of two spaces are linked by the new matrix elements, rather than the Minkowski metric. Our calculations show that the law of direct and crossed multiplication between $q$-vectors is given by $H^{\mu\nu}=\left(\begin{array}[]{cccc}1&\imath&\imath&\imath\\ -\imath&-1&1&-1\\ -\imath&-1&-1&1\\ -\imath&1&-1&-1\\ \end{array}\right)$, where
@@
We also introduce a new tensor, denoted as $n_{k}^{\ ij}$, derived from the behavior of quaternions. This tensor has the following properties: it is antisymmetric concerning the exchange of indices, and can be used to raise and lower indices; it is spatial, although it does not involve counting, and it includes a temporal index. The $n_{k}^{\ ij}$-symbol is used when its indices are all on the same row. When all indices are in a row, it resembles the Levi-Civita symbol. We will identify the location of the temporal index later.
@@
where $\mathcal{E}_{\mu\nu}$ is given by Eq.I. The equation of motion describes the action of the covariant derivative on the corresponding quaternion variables and vice versa. Due to the incompleteness of the derivative effects, to obtain the equation of motion, we must define a new commutator as follows: $[\ ,\ ]_{\triangleright}:\ \mathsf{Q}\times\mathsf{Q}\rightarrow\mathsf{Q}$ which is the same as $[\mathtt{B},\mathtt{C}]_{\triangleright}=(\eta^{\mu\nu}[\mathtt{B}_{\mu},\mathtt{C}_{\nu}]_{\ast},[\mathtt{B},\mathtt{C}]^{\vee}_{\ast}+[\mathtt{B}\mathtt{B},\mathtt{C}\mathtt{C}]^{\wedge}_{\ast})$. Due to the electrodynamics in Minkowski spacetime, the commutator of the motion has the following form (under the matrix $\tilde{H}$):
@@
To reach the final goal (Eq.26), the relation $\mathtt{A}_{s;s}-\mathtt{A}_{0;0}=\mathtt{A}_{s}^{\ ;\zeta}g_{s\zeta}-\mathtt{A}_{0}^{\ ;\zeta}g_{0\zeta}$ must be established, and this is a precondition. With the above gauge, the continuity equation will be: $\mathsf{j}_{0}^{\ ;0}+\mathsf{j}_{k}^{\ ;k}=0$.
@@
III Conclusions
We demonstrate that proving gravitational issues, such as dark matter, Einstein’s equation can’t take any additional terms. Modifying Einstein’s equation is possible only by vanishing covariant derivative terms.
IV Acknowledgments
V Geometry
@@
The new notation $\mathcal{F}_{\mu\nu}$ (Eq.I) serves to visually distinguish this variable from that of the previous section, $\mathcal{E}_{\mu\nu}$. As a result, all previous relationships and their arrangements remain intact with this transformation. Thus, the relation in 18 will be the equation of motion for the new fields $\mathcal{F}_{\mu\nu}$. By entering the quaternion current density and separating the main equations into real and imaginary components, we have
@@
The connection between the geometric tensor dynamics and sources is clear. Comparing the results from the previous section, it is evident that the properties of the gauge field strength tensor apply to both components of Riemann’s curvature tensor. These performances emphasize that the Christoffel connections have the gauge property. Equations 27 and V give the source less equations:
@@
Eq.30 represents the Bianchi identities and can be derived from $\hat{G}^{\mu\nu}=\frac{1}{2}\varepsilon^{\mu\nu\alpha\beta}\hat{R}_{\alpha\beta}$. For the Bianchi identities, the sub-components of the released space behave independently. But, in the second relation, which is independent of the source, the behavior of the second Bianichi identity is intertwined with the basis of the subcomponents. The remaining two relations of the basic equation of motion (Eq.V) are as follows:
@@
On the other hand, if we use the equation of motion (Eq.V)
@@
is like the famous case $\mathtt{C}_{\lambda\nu}=-\Lambda g_{\lambda\nu}$. It turns out that $\mathtt{C}_{\lambda\nu}=f_{\lambda\nu}(\mathbf{g})$ (for example $(e^{\kappa\mathbf{g}})_{\lambda\nu}$, $\mathbf{g}$ is the matrix for the metric). $\mathtt{C}_{\lambda\nu}$, is a geometrical quantity, not matter. It can be considered general background matter, which can be absorbed by the tensor $\mathtt{T}_{\lambda\nu}$ and become zero.
@@
In the final equation, the characters of geometry and matter are distinguished, from one another. The fact that in Einstein’s theory of gravitation, Einstein’s term (in agreement with the covariant derivative) is not the only way to choose $\mathtt{C}_{\lambda\nu}$ is another result of this mechanism. We must also consider constants that indicate the limitations of spacetime to solve the equation of gravity. It is a question of finding the constant from equation 36 because it is not a unique quantity. Changing any principal constant in equation 36 (agreement with covariant derivative) leads us into a new universe. Among the possible options for the constant $\mathtt{C}_{\lambda\nu}$, equation 36 makes a perturbation theorem. What can be obtained, however, is the non-uniqueness of Einstein’s famous theorem.
VI Discussion

### v3

@@ Abstract
We treat the Christoffel coefficients as operators and introduce new mappings for quaternionic products to connect with the theory of electrodynamics in general spacetime. By utilizing the directional operator of the covariant derivative, we generalize the quaternionic mechanism to the theory of electrodynamics. We demonstrate that the Einstein equation permits the selection of a constant term that is consistent with the covariant derivative.
@@ I Introduction
Many attempts were made to develop a geometric approach to electrodynamics when general relativity successfully explained the universe. Unfortunately, most of these efforts were unsuccessful overduin ; ruiz ; wes ; cabral ; goenner ; lobo ; wess ; gross ; ponce ; williams ; wesso . Our current goal is to pursue this objective in a significantly different manner. We aim to incorporate the theory of gravity into the framework of electrodynamics. This integration may enhance our understanding of quantum gravity, a concept that physicists widely seek to realize. To validate the results of this work, we can analyze the components of Riemann’s curvature tensor. The quest for a deeper understanding of quantum gravity has prompted us to propose a modification in the dimensions of spacetime herbert ; kiefer ; gianluca ; gianlu ; gianluc ; carlo ; ashteka ; ashtekar . In addressing this challenge, we encounter local issues related to quaternionic production, particularly regarding the contraction of indices in tangent and cotangent spaces. However, the fundamental outcome remains consistent, regardless of the chosen space. These new perspectives will be instrumental in advancing our understanding of gravity theory. We examine two structures: the spacetime metric and the quaternionic metric, which involves quaternionic production under a new metric. As will be demonstrated, it is possible to transition from electrodynamics to gravity using quaternions. However, there is a fundamental difference between the two theories. Electrodynamics is a comprehensive theory involving two components, described by the strength tensor in four-dimensional space. When extending electrodynamics to a theory of gravity using quaternions, we must consider four components in a four-dimensional spacetime. This leads to the idea of component separation. The alteration in the structure of these components results in equations of motion, ultimately culminating in Bianchi’s identity weber ; misner ; wald ; weinberg ; invarno ; malcolm . We will show that space (geometry) has a physical entity. Therefore, equations of motion will have to be determined, which go beyond mathematical relations to the properties of the geometry of space. When space takes on a physical meaning, it must obey the laws of physics, and ”space characters behavior must be affected by the equations of motion”. One consequence is that Einstein’s equation for gravity is merely one of the equations that determine geometry.
@@
Clarification- It is essential to clarify definitions and designations: the vector space is represented by $\varepsilon$ and its dual vector space is denoted by $\varepsilon^{\star}$. They are ordered and denoted as $\varepsilon^{\star}\otimes\varepsilon$, and other arrangements are not permitted. The notation $\mathsf{u}^{\alpha\beta\cdots}_{\mu\nu\cdots}=u^{\alpha\beta\cdots}_{\mu\nu\cdots}\,(\hat{e}^{\mu}\hat{e}^{\nu}\cdots)\otimes(\hat{e}_{\alpha}\hat{e}_{\beta}\cdots)$ indicates a multicomponent member. The notation is revised as follows: the components $\mathsf{u}^{\alpha\beta}_{\mu\nu}$ will now be expressed as $u^{\alpha\beta}_{\mu\nu}$. The interaction of indices occurs only from a vector space to the dual space and vice versa; in any case, it is from a vector space into the dual space.
@@
The symbol "$\mathbf{g}"$ is the general metric element and comprises $\varepsilon\times\varepsilon$ or $\varepsilon^{\star}\times\varepsilon^{\star}$. Equations 1 and 2 show that the metric (with a repeated index) does not obey the rules of the operator, and it is clear that any other interpretation of equations 1 and 2 would be wrong for the action of the metric. It is demonstrated that the metric effect does not influence the calculation of the components. Following the definition of "$\mathsf{ast}"$ mapping, the operator $\hat{\Gamma}_{\kappa}$ is represented as follows;
@@
The above statements and results become easier to understand with the new representation because the $\hat{\Gamma}_{\zeta}$ operator yields "$\hat{\Gamma}_{\kappa}=\Sigma_{\alpha,\beta}\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|\equiv\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|"$. The "$\mathsf{ast}"$ mapping becomes more formal when referring to the bracket representations. Moreover, the representation corresponding to the bracket extends our possibilities for future calculations. The $\alpha$ and $\beta$ are dummy indices in the $\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|$-operator. They are used, among other things, for the gamma operator effect;
@@
During this process, the box is aligned with the alpha vector. That is, the box takes the alpha index. Hence,
@@
The relationships between the quaternion generators $\hat{q}_{1}=I$, $\hat{q}_{2}=J$, and $\hat{q}_{3}=K$ are defined within a Euclidean metric and adhere to the following properties: $\hat{q}_{i}\cdot\hat{q}_{j}=\delta_{ij}$ and $\hat{q}_{i}\times\hat{q}_{j}=\Sigma_{k}\epsilon_{ijk}\hat{q}_{k}$. If we include $\hat{q}_{0}=\Im$ (as the identity), we obtain a quaternion represented by the Lorentz quantities (parameters of the Lorentz index). These are coupled and are termed “$q$ vectors.” Lorentz quantities, which carry the Lorentz index, are represented using quaternions. We have two representations of the Lorentz index: the four-vector and the q-vector. We will use the notation $\mathtt{A},\ \mathtt{B},\cdots$ to denote the quaternions. Quaternions are typically constructed from four vectors, denoted as $A_{\mu}=(A_{0},\mathbf{A})$. They can be expressed as $\mathtt{A}=iA_{0}\Im+A_{k}\hat{q}^{k}\equiv(\imath\mathtt{A}_{0},\mathtt{A})_{q}$, where the $q$-variable signifies the expression $(\imath\mathtt{A}_{0},\mathtt{A}_{k}\hat{q}^{k})$. Bold letters denote the spatial components of the four vectors. The four-vector framework is based on the spacetime metric, while the Euclidean metric pertains to the quaternion representation. The configuration significantly affects quaternionic products. We also analyze the Lorentz index, which connects physical parameters to the quaternion generator. We can generalize the range of coupled quantities from vectors to tensors. For the $q$-vectors $\hat{\Gamma}$ and $\mathtt{C}$, the extended Grassman quaternion multiplication """"$\triangleright$"" ▷ " (with $\ast$-product), is defined as follows andre ; walker ; kansu ; stefano
@@
where "$e"$ is the coupling coefficient constant with the value of one. As mentioned earlier, $\hat{\Gamma}_{\mu}$ contains two coupling indices. Due to the operators $\hat{\Gamma}$’s and the $\ast$-product $\Delta_{\mu}$ is a semi-complete operator. Equation 7 with the $\ast$-product has the expected effects
@@
One can observe that $g^{\alpha\beta}_{\ \ ;\beta}=0$, and $\mathtt{B}^{\lambda}_{\ ;\lambda}=\mathtt{B}_{\lambda}^{\ ;\lambda}$. The covariant derivative is not complete. The incompleteness of the derivative operator is evident from equations I and 5.
$\mu$ and $\zeta$ are free indices in $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$. This means that $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$ ends up being a second rank tensor; In other words, $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}\propto\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}$ actually has the two free indices $\mu$ and $\zeta$ that can be considered as free indices; $\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\equiv\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\ |\mu,\zeta>$. Equation 3, and our calculations in the bracket notation show that,
@@
The free indices following the transposition are visible. In this way, $(\Gamma^{\zeta}_{\mu\lambda}\Gamma^{\kappa}_{\nu\zeta}-\Gamma^{\zeta}_{\nu\lambda}\Gamma^{\kappa}_{\mu\zeta})(|\kappa><\lambda|)$ as an operator contains the two free indices $\kappa$ and $\lambda$. It becomes a simplicial $[\hat{\Gamma}_{\mu},\hat{\Gamma}_{\nu}]^{\kappa}_{\ast\ \lambda}$ which has free $\lambda$ and $\kappa$ indices. Next, we follow the computation of $[\Delta_{\alpha},\Delta_{\beta}]_{\ast}$. We compute the case without torsion,
@@ II Electrodynamics
In this section, we present a mechanism by which the components of gauge fields are coupled with the generators of quaternions ($\mathbf{A}\rightarrow\mathtt{A}$). A connection has been established between the components of the covariant derivative and the quaternion generators. The quaternion variables obtained in this way help us to reconstruct the strength tensor of the fields. Due to the semi-completeness of the covariant derivative operator, we will introduce a new quaternion commutator to obtain the equation of motion and conclude the theory.
@@
We show that Maxwell’s equations are accessible with augmented quaternions. The given content confirms the above definitions of the $q$-vectors. We emphasize that the quaternion mapping creates two separate spaces: the tangent and the cotangent $q$-vectors. The measures of the spaces are related by the new matrix elements, rather than by the Minkowski metric. Our calculations show that the law of direct and crossed multiplication between $q$-vectors is given by $H^{\mu\nu}=\left(\begin{array}[]{cccc}1&\imath&\imath&\imath\\ -\imath&-1&1&-1\\ -\imath&-1&-1&1\\ -\imath&1&-1&-1\\ \end{array}\right)$, where
@@
We also introduce a new tensor, denoted as $n_{k}^{\ ij}=g_{ks}\epsilon^{sij}$, derived from the behavior of quaternions. This tensor has the following properties: it is antisymmetric concerning the exchange of indices and can be employed to raise and lower indices; it is spatial, although it does not involve counting, and it includes a temporal index. Symbole $n_{k}^{\ ij}$ is used when its indices are on the same row. When all indices are in a row, it resembles the Levi-Civita symbol. We will identify the location of the temporal index later.
@@
where $\mathcal{E}_{\mu\nu}$ is given by Eq.I. The equation of motion describes the action of the covariant derivative on quaternion variables and vice versa. Due to the incompleteness of the derivative effects, to obtain the equation of motion, we must define a new commutator as follows: $[\ ,\ ]_{\triangleright}:\ \mathsf{Q}\times\mathsf{Q}\rightarrow\mathsf{Q}$ which is the same as $[\mathtt{B},\mathtt{C}]_{\triangleright}=(\eta^{\mu\nu}[\mathtt{B}_{\mu},\mathtt{C}_{\nu}]_{\ast},[\mathtt{B},\mathtt{C}]^{\vee}_{\ast}+[\mathtt{B}\mathtt{B},\mathtt{C}\mathtt{C}]^{\wedge}_{\ast})$. Due to the electrodynamics in Minkowski spacetime, the commutator of the motion has the following form (under the matrix $\tilde{H}$):
@@
To reach the final goal (Eq.26), the relation $\mathtt{A}^{\kappa}_{\ ;\kappa}=0$ must be established, and this is a precondition. With the above gauge, the continuity equation will be: $\mathsf{j}^{\kappa}_{\ ;\kappa}=0$.
@@
III Geometry
@@
The new notation $\mathcal{F}_{\mu\nu}$ (Eq.I) serves to visually distinguish this variable from that of the previous section, $\mathcal{E}_{\mu\nu}$. As a result, all previous relationships and their arrangements remain intact with this transformation. Thus, the equation of motion for the new fields $\mathcal{F}_{\mu\nu}$ will be found in the relation in 18. By entering the quaternion current density and separating the main equations into real and imaginary components, we have
@@
There is a clear connection between the sources and dynamics of geometric tensors. Considering the results from the previous section, it is evident that the properties of the gauge field strength tensor apply to both pair components of the Riemann curvature tensor. These performances emphasize that the Christoffel connections have the gauge property. Equations 27 and III give the source less equations:
@@
Eq.30 represents the Bianchi identities and can be derived from $\hat{G}^{\mu\nu}=\frac{1}{2}\varepsilon^{\mu\nu\alpha\beta}\hat{R}_{\alpha\beta}$. For the Bianchi identities, the sub-components of the released space behave independently. But, in the second relation, which is independent of the source, the behavior of the second Bianichi identity is intertwined with the basis of the subcomponents. The remaining two relations of the basic equation of motion (Eq.III) are as follows:
@@
On the other hand, if we use the equation of motion (Eq.III)
@@
is like the famous case $\mathtt{C}_{\lambda\nu}=-\Lambda g_{\lambda\nu}$. It turns out that $\mathtt{C}_{\lambda\nu}=f_{\lambda\nu}(\mathbf{g})$ (for example $(e^{\kappa\mathbf{g}})_{\lambda\nu}$, $\mathbf{g}$ is the matrix for the metric). $\mathtt{C}_{\lambda\nu}$, is a geometrical quantity, not matter. The tensor $\mathtt{T}_{\lambda\nu}$ absorbs the general background matter influence, and the effect becomes zero.
@@
In the final equation, the characteristics of geometry and matter are distinguished. The fact that in Einstein’s theory of gravitation, Einstein’s term (in agreement with the covariant derivative) is not the only way to choose $\mathtt{C}_{\lambda\nu}$ is another result of this mechanism. We must also consider the constants that indicate the limitations of spacetime when solving the equation of gravity. It is a question of finding the constant from equation 36 because it is not a unique quantity. Changing any principal constant in equation 36 (agreement with covariant derivative) leads us into a new universe. Among the possible options for the constant $\mathtt{C}_{\lambda\nu}$, equation 36 makes a perturbation theorem. What can be obtained, however, is the non-uniqueness of Einstein’s famous theorem.
IV Conclusions
The results of this article may provide solutions to gravitational issues such as dark matter and be forbidden on Einstein’s equation to take additional terms. It is possible to modify Einstein’s equation by adding terms that have vanished covariant derivatives.
V Discussion